OpenStudy (anonymous):

What is the value of y = sin (θ) when θ = pie /2 ?

OpenStudy (anonymous):

@bibby

OpenStudy (anonymous):

The value is 1

OpenStudy (jdoe0001):

\(\bf y=sin(\theta)\qquad \theta=\cfrac{pie}{2}\implies half-pie\\ \quad \\ y=sin(half-pie)\implies y= \textit{whipped cream}\\ \quad \\ \textit{unless you meant }\qquad \theta=\cfrac{\pi}{2}\qquad then\\ \quad \\ y=sin\left(\frac{\pi}{2}\right)\) check your Unit Circle